# Efficient Exploration and Discriminative World Model Learning with an Object-Centric Abstraction

### v1

@@ Appendix E Baseline Methods
Specifically, we use Dreamer-v3 as a performant model-based baseline (Hafner et al.,, 2023), and PPO as a model free baseline (Schulman et al.,, 2017; Petrenko et al.,, 2020). For exploration-reward transfer experiments, we use a disagreement-based model for Dreamer intrinsic reward (Sekar et al.,, 2020), as well as using the same count-based reward to train all models.
@@ F.1 Extended Main Results
All runs, extended and main, are run with a minimum of 3 independent seeds.
To get the performance of the various algorithms in the low level MDP as we plotted them in Figure 4(b), we extracted them from Figure 13 in Henaff et al., (2022), using the WebPlotDigitizer app (https://apps.automeris.io/wpd/).
@@ F.1.2 Dreamer
F.1.3 Non-Parametric Model
We plot comprehensively the non-parametric model results in Figure 21.
@@ F.2 Low Level Learning
While this can be implemented in many ways, in practice we use a goal conditioned RL set-up, where a single model free policy network takes in both the low level state and the behaviour embedding to produce a primitive action: $\pi:\mathcal{S}\times\mathcal{B}\rightarrow\mathcal{A}$. The low level state is the environment’s raw observations while the behaviour embedding is the same as the one given to the abstract level model. This is implemented as a wrapper over the original environment which provides the policy with a +1 reward if $(\alpha^{(i)},\xi^{\prime})\in X_{t+1}$, and -1 otherwise. We use APPO as the policy optimization algorithm (Petrenko et al.,, 2020). The point of this is merely to show that we can learn the set of behaviours neurally and building an Ab-MDP with them still works.
@@ F.3.2 Model fitting
F.3.3 Semantic Embeddings
F.3.4 GloVe Embedding
@@
We investigate the impact of having the object identity embeddings be semantically meaningful versus random one-hot embedding in MiniHack. By default, all experiments in the main text take as input object identity embedding using GloVe Pennington et al., (2014) embedding, which are kept frozen throughout training. We ablate this by initializing the object identity embedding matrix to be randomly initialized from a standard Gaussian, $\mathcal{N}(0,1)$, and allow them to be updated by gradients.
@@ F.4 Model Updates
We show effects of number of model update steps during data collection in Figure 26,27,&28.
@@ G.0.1 Our model
G.0.2 Dreamer-v3
We describe the default parameter used to run the Dreamer-v3 baseline. We use a PyTorch implementation of Dreamer: https://github.com/NM512/dreamerv3-torch, which has similar training curves as reported initially in Dreamer-v3. Unless otherwise mentioned, we use the “XL” world model setting with the highest amount of training iteration, which had the best reported sample efficiency in Hafner et al., (2023), and was confirmed by our preliminary experiments. We start with parameters as faithful to the reported ones in Hafner et al., (2023) as possible as it was found to work across a wide range of settings, then sweep over a subset of hyperparameters and report the best ones (which we use) in Table 5.
We perform parameter sweep over a subset and report the best ones (which we use) in Otherwise we keep the parameters as faithful to the reported ones in Hafner et al., (2023) as possible as it was found to work across a wide range of settings.
@@ G.0.2 Dreamer-v3
G.0.3 PPO
We use the Sample-Factory implementation of PPO (Petrenko et al.,, 2020). We run a parameter sweep over a subset and report the best parameters which we use in Table 6.

### v3

@@ Appendix E Baseline Methods
Specifically, we use Dreamer-v3 as a performant model-based baseline (Hafner et al., 2023), MuZero as another model-based baseline (Schrittwieser et al., 2020; Werner Duvaud, 2019), and PPO as a model free baseline (Schulman et al., 2017; Petrenko et al., 2020). For exploration-reward transfer experiments, we use a disagreement-based model for Dreamer intrinsic reward (Sekar et al., 2020), as well as using the same count-based reward to train all models.
Finally, we also compare against other methods that uses a structured MDP, such as the attribute planner in Zhang et al. (2018), and NCS in Chang et al. (2023). These results can be found in Section F.1.3.
We outline the main differences between the different model-learning methods in Table 5.
State Representation Function Approx Planning
Dreamer-V3 (Hafner et al., 2023) Feature vector $\phi$ Neural Net (generative) Model free (in imagination)
MuZero (Schrittwieser et al., 2020) Feature vector $\phi$ Neural Net (generative) MCTS
DAC-MDP (Shrestha et al., 2020) Feature vector $\phi$ k-NN Value iteration (on k-NN “cores”)
Attribute Planner (Zhang et al., 2018) Tabular No approximation Dijkstra
DLP-EIT (Haramati et al., 2024) Sets of vectors Neural Net (generative) Model free
NCS (Chang et al., 2023) Sets of $(\alpha,\xi)$ vectors K-means (on just attribute vectors $\xi$) Greedy
MEAD Sets of $(\alpha,\xi)$ vectors Neural net (discriminative) Dijkstra
Ablation Sets of $(\alpha,\xi)$ vectors K-means (on just attribute vectors $\xi$) Dijkstras
Table 5: Model-based methods
@@ F.1 Extended Main Results
All runs, extended and main, are run with a minimum of 3 independent seeds. We outline the specific number of seeds for each method and each main experiment in Table 6.
Experiment Group Method N Seeds
Learning from scratch, 2D crafting games (Figure 4(a)) MEAD 4
Dreamer-V3 3
PPO 3
Learning from scratch, Minihack skill games (Figure 4(b)) MEAD 5
MuZero 3
Dreamer-V3 4
PPO 3
Transfer from Sandbox pre-training (Figure 5(a)) MEAD 3
Dreamer-V3 (count / disagreement / reward) 3 / 3 / 3
Transfer from Freeze-Random pretraining (Figure 5(b)) MEAD 3
Dreamer-V3 (count / disagreement / reward) 3 / 3 / 3
Compositional Environment (Figure 5(c)) MEAD (from scratch / pretrained) 3 / 3
Dreamer-V3 (from scratch / pretrained) 3 / 3
Table 6: Number of seeds used in our main results.
To get the performance of the various algorithms in the low level MDP as we plotted them in Figure 4(b), we extracted them from Figure 13 in Henaff et al. (2022), using the WebPlotDigitizer app (https://apps.automeris.io/wpd/).
@@ F.1.2 Dreamer
F.1.3 Other Model Learning Approaches
We plot comprehensively the non-parametric model results in Figure 21. This model uses a table to keep track of transition successful probabilities between abstract states, as used by the attribute planner in Zhang et al. (2018). We show direct comparison between methods in Figure 17.
@@ F.1.3 Other Model Learning Approaches
We further compare against an approach using K-means clustering on the attributes, and labeling edges (between the k-means “cores”) with the last observed action that successfully led to a transition. This way of model-learning along with greedy planning was used by the NCS (neural constraint satisfaction) method of Chang et al. (2023). Specifically NCS uses k-means clustering with greedy planning, which we found to not work for any of our environments. A hybrid variant using k-means clustering with Dijkstra planning works in MiniHack (albeit less well than MEAD), and fails completely in 2D crafting.
(a) Training from scratch in 2D crafting env.
(b) Training from scratch in MiniHack.
Figure 23: Performance of the k-means clustering on attributes method as used in NCS (Chang et al., 2023). NCS specifically uses k-means clustering with greedy planning, although we also test a variant using k-means clustering with Dijkstra planning. Legends specify the function approximation and planning methods used.
@@ F.2.1 Learning Low Level Policies
While this can be implemented in many ways, in practice we use a goal conditioned RL set-up, where a single model free policy network takes in both the low level state and the behaviour embedding to produce a primitive action: $\pi:\mathcal{S}\times\mathcal{B}\rightarrow\mathcal{A}$. The low level state is the environment’s raw observations while the behaviour embedding is the same as the one given to the abstract level model. This is implemented as a wrapper over the original environment which provides the policy with a +1 reward if $(\alpha^{(i)},\xi^{\prime})\in X_{t+1}$, and -1 otherwise. We use APPO as the policy optimization algorithm (Petrenko et al., 2020). The point of this is merely to show that we can learn the set of behaviours neurally and building an Ab-MDP with them still works.
@@
F.2.2 Learning Object-Centric Map
We train an object centric map which takes the low level observation and predict the abstract state: $\hat{X_{t}}=M_{\theta}(S_{t-1},A_{t},S_{t})$. The encoder map $M_{\theta}$ uses a convolutional encoder for MiniHack low level observations (specifically, the low level encoder from Henaff et al. (2022)), and learned embeddings for each discrete actions. After concatenating all encoded inputs, the latent is projected through two 2-layer feedforward nets, one producing a categorical distribution over item identities, and another a categorical distribution over item attributes. The output of this map are two tensors each with shape (item set size $\times$ num categories). We can train this via a categorical cross-entropy loss to predict the correct identity and attribute categories for all items.
We collect a dataset of 100k transitions using a random policy in the MiniHack Freeze-Horn environment, and use the hand-crafted mapping to provide ground truth abstract state labels. We train using Adams optimizer with batch size 4096 and learning rate 3e-4.
We investigate the effect of training with different sub-sets of the data on both the encoder accuracy and the subsequent planning performance in Figure 25.
Figure 25: Effect of training on different proportions of the 100k dataset on encoder accuracy (left) and MEAD performance (right) in Minihack-Freeze-Horn.
@@ F.3.2 Model fitting
F.3.3 GloVe Semantic Embedding
@@
We investigate the impact of having the object identity embeddings be semantically meaningful versus random one-hot embedding in MiniHack. By default, all experiments in the main text take as input object identity embedding using GloVe Pennington et al. (2014) embedding, which are kept frozen throughout training. We ablate this by initializing the object identity embedding matrix to be randomly initialized from a standard Gaussian, $\mathcal{N}(0,1)$, and allow them to be updated by gradients.
@@ F.4 Model Updates
We show effects of number of model update steps during data collection in Figure 28, 29,&31.
@@
Figure 30: Ablating just the effect of model optimization steps in MiniHack-Freeze-Wand env. Model weights are reset after each data collection step (of 2.5k frames), then trained for the specified number of steps before the next data collection step. With sufficient optimization steps, high model accuracy and planning performance is achieved.
@@ G.1 Our model
G.2 Dreamer-v3
We describe the default parameter used to run the Dreamer-v3 baseline. We use a PyTorch implementation of Dreamer: https://github.com/NM512/dreamerv3-torch, which has similar training curves as reported initially in Dreamer-v3. Unless otherwise mentioned, we use the “XL” world model setting with the highest amount of training iteration, which had the best reported sample efficiency in Hafner et al. (2023), and was confirmed by our preliminary experiments (Figure 35). We start with parameters as faithful to the reported ones in Hafner et al. (2023) as possible as it was found to work across a wide range of settings, then sweep over a subset of hyperparameters and report the best ones (which we use) in Table 8.
We perform parameter sweep over a subset and report the best ones (which we use) in Otherwise we keep the parameters as faithful to the reported ones in Hafner et al. (2023) as possible as it was found to work across a wide range of settings.
@@ G.2 Dreamer-v3
(a) Training from scratch (no pre-training)
(b) Fine-tuning after reward-based pre-training
Figure 35: Effect of Dreamer-v3 training ratio (amount of training to data collection) on performance, for training from scratch in MiniHack-Freeze-Wand (left), and for fine-tuning in MiniHack-Levitate-Boots from reward-based pretraining in MiniHack-Freeze-Random.
G.3 MuZero
We use the MuZero code base from Werner Duvaud (2019). We sweep over learning rates $1\times 10^{-5}$, $1\times 10^{-4}$, $1\times 10^{-3}$}, encoding size {32, 64, 128, 256, 512}, number of simulations {30, 90, 100, 180}, using prioritized experience replay (or not), batch size {128, 2048}, TD steps {4, 10, 20, 50, 100}, and the delay in self-play simulations (allowing for more training steps in between data collections) of {0, 2, 8}. We use the best hyperparameters and report them in Table 9.
Hyperparameter Value
Learning rate init $1\times 10^{-4}$
Encoding size 512
Batch size 2048
Num simulations 100
TD steps 20
Prioritized experience replay False
Table 9: MuZero parameters
We observe a small effect of different self play delay lengths performing better in different games. We choose a delay of 0 as it performed the best in most games. Nevertheless, we report the performance of different self play delay lengths in Figure 36 for completeness.
Figure 36: Muzero runs in the Minihack environments, with different self-play delay lengths
G.4 PPO
We use the Sample-Factory implementation of PPO (Petrenko et al., 2020). We run a parameter sweep over a subset and report the best parameters which we use in Table 10.